# Ultrafast broadband soliton microcomb laser

### v1

@@ Abstract
Optical frequency combs, defined by their ultrafast timescale and profound coherence, have become a foundational technology in science and technology. Their chip-scale miniaturization offers great promise for refining a wide range of applications from metrology to advanced computing. Existing microcomb technologies, however, are constrained by fundamental limitations inherent to their mode-locking mechanisms. Specifically, the carrier dynamics of saturable absorption in semiconductor mode-locked lasers and the external laser pumping required for dissipative Kerr soliton generation lead to degraded comb performance and significant operational complexity. Here, we introduce an on-chip mode-locking approach that resolves these challenges. By leveraging the intriguing property of a thin-film lithium niobate photonic circuit integrated with semiconductor optical gain, the chip-scale soliton laser directly emits background-free, ultrashort soliton pulses with a 3-dB bandwidth exceeding 3.4 THz (<<90 fs pulse width) and an ultra-narrow comb linewidth down to 53 Hz. Our approach dramatically simplifies bright-soliton generation, featuring fully electric pumping, stable turn-key operation, near-unity optical efficiency, superior long-term stability, and remarkably low soliton generation threshold (1 V, 75 mA) that could be powered with a standard alkaline battery. The demonstrated soliton laser represents a paradigm shift in chip-scale frequency comb technology, which is expected to have profound impacts across a wide range of microcomb applications.
@@ Introduction
Optical frequency combs (OFCs) – coherent light sources comprised of equidistant comb lines with phases locked with each other – form the foundation for many important research areas such as metrology, sensing, clock, and photonic signal processing 1, 2, 3. Producing OFCs on chip, so-called soliton microcombs, is essential for reducing the size, weight, and power consumption of OFC systems, whose development in recent years 4, 5, 6, 7, 8 has significantly advanced numerous applications including communication 9, spectroscopy 10, ranging 11, 12, 13, optical and microwave synthesis 14, 15, 16, and computing 17, 18. Despite these great successes, current soliton microcomb technologies have remained predominantly confined to research laboratory demonstrations, primarily due to the fundamental challenges lying in their very nature of comb generation and mode-locking mechanisms that hinder their transition from laboratory settings to broad consumer applications.
To date, soliton microcombs are produced primarily by two dominant approaches 4, 7, 8, either via externally driven parametric processes exemplified by dissipative Kerr solitons (DKS), or via internally driven parametric processes represented by semiconductor mode-locked lasers (SMLLs). An SMLL exploits optical gain in conjunction with saturable absorption in an active semiconductor laser cavity for soliton comb generation (Fig. 1a). The soliton pulses are sustained by the optical gain, with unutilized energy stored in the upper energy levels of the gain medium for amplifying subsequent pulses, resulting in background-free soliton pulses with 100% utilization of optical power for pulse generation. As the nonlinear process for soliton mode locking is entirely internal to the laser cavity, it is fully electrically turnkey and thus represents a convenient path for producing soliton microcomb. Unfortunately, SMLLs face fundamental challenges in scaling toward a short pulse width and/or a high repetition rate in that the complex carrier dynamics associated with saturable absorption and gain recovery seriously limit the modulation speed and destabilize the formation of ultrashort pulses. Consequently, SMLLs developed so far show only a limited pulse width of sub-picoseconds and a repetition rate in the order of $\sim$100 GHz  7, 8.
Figure 1: Concept and implementation of soliton microcomb laser. (a) Schematic illustration of different device structures and mode-locking mechanisms in a semiconductor mode-locked laser (SMLL), a dissipative Kerr soliton (DKS) generator, and our soliton microcomb laser. A SMLL employs saturable absorption for mode locking that fundamentally limits pulse bandwidth and speed. DKS relies on external pumping of a Kerr resonator to access broadband parametric gain, yet suffers from low conversion efficiency and device thermal dynamics that complicates soliton access. Our soliton microcomb laser utilizes a photorefractive-stabilized microresonator inside a III-V/TFLN external laser cavity for mode locking, which enables spontaneous generation of broadband soliton pulses with direct electrical pumping. (b) Photograph (top) and schematic (bottom) of the hybrid integrated laser architecture. (c) Optical micrographs detailing the RSOA-TFLN coupling interface, embedded high-Q microresonator, and broadband Sagnac mirror. (d) Photograph showing the device footprint in comparison with an one-cent coin.
DKS systems exploit an external laser to pump a passive optical Kerr microresonator for soliton comb generation (Fig. 1a). The dramatically enhanced parametric process inside the high-Q resonator results in a broad comb bandwidth and exceptional coherence property, which have garnered significant interest in the past decade. However, the nature of external driving with a pump laser also introduces fundamental challenges to DKS generation. On one hand, it requires sophisticated frequency locking of the pump laser to the high-Q optical resonance for proper operation, which is further complicated by the significant thermo-optic nonlinearity that perturbs the soliton initiation. On the other hand, it suffers from low efficiency of converting the continuous-wave (CW) pump laser to the short soliton pulses due to their limited temporal overlap for energy exchange, leading to a significant pump-wave background underlying the produced soliton pulses. Furthermore, the comb linewidth is primarily confined by that of the pump laser 19, 20, 21. Although the efficiency can be improved by pulse pumping or pump power recycling 22, 23, 24, 25, they require delicate synchronization in time or resonance frequency and the difficulty of soliton initialization remains the same. The challenge of soliton triggering can be resolved with certain self-starting techniques 26, 27, 28, 29, 30, while low conversion efficiency persists. Certain “active resonator” approaches 31 were recently introduced to address some of these issues by integrating with erbium-doped fiber amplifier (EDFA) gain 32, 33, 34 or quantum cascade laser (QCL) gain 35, 31, which, however, result in limited comb bandwidths and face challenge in on-chip integration due to the requirement of bulky off-chip EDFA 32, 33, 34 or external laser driving 31. To date, a DKS source free from all these challenges remains elusive.
Here, we demonstrate a soliton microcomb laser that unites elegantly the advantages of SMLL for convenient turn-key operation with 100% utilization of optical power for pulse generation, and DKS for broad comb bandwidth with high coherence, while free from all the challenges they encounter. The chip-scale laser directly emits broadband soliton pulses with a 3-dB bandwidth of >>3.4 THz in the telecom band and a corresponding ultrashort pulse width of <<90 fs, and an ultrahigh repetition frequency up to 0.8–3 THz, significantly beyond the reach of conventional SMLLs. Moreover, the emitted soliton comb exhibits an ultra-narrow linewidth down to only 53 Hz, the smallest linewidth ever demonstrated for on-chip soliton microcombs including DKSs and SMLLs 7, 8, 36. The laser is electrically pumped with stable turn-key operation and near-unity optical efficiency, same as SMLL. The laser exhibits remarkably low soliton generation thresholds with a driving voltage and current of 1 V and 75 mA that is orders of magnitude lower than DKSs. The demonstrated soliton comb laser represents a paradigm change in soliton microcomb generation, which is expected to have a profound impact on broad OFC applications in communication, sensing, metrology, computing, and beyond.
@@
Figure 2: Performance of the soliton microcomb laser. (a) Optical (left) and RF (right) spectra of laser output at different RSOA driving currents. Dashed lines represent $sech^{2}$ fits. Each state was obtained by optimizing the relative position between the RSOA and TFLN chips to compensate for thermal-induced mechanical drift. (b) FROG traces for the 1.2-THz (top), 1.6-THz (middle), and 0.4-THz (bottom) soliton states at RSOA driving currents of 75, 120, and 151 mA, respectively. The center frequency ($\Delta\nu=0$) of the FROG traces is 386.83THz (corresponding wavelength of 775 nm). For a better comparison, the figures also show frequency and time scales normalized, respectively, by the 200-GHz FSR and 5-ps roundtrip time ($T_{0}$) of the mode-locking resonator. The relatively indistinct pulse temporal boundaries in the FROG traces result simply from limited gain bandwidth of the EDFA used to boost the soliton pulses before they are recorded by FROG (Fig. 2d). This artifact is confirmed by performing the same FROG on similar comb states produced via conventional DKS excitation in a passive microresonator. Details are provided in SI. (c) Comparison with advanced SMLLs on optical bandwidth and repetition rate. All data are listed in Table S1 of SI. (d) Schematic of experimental setup. LDC: laser diode controller; DCU: dispersion compensation unit; PM: power meter; FROG: frequency-resolved optical gating; OSA and ESA: optical and electrical spectrum analyzer.
@@
Our laser integrates an InP‑based reflective semiconductor optical amplifier (RSOA) as the gain medium with a thin-film lithium-niobate (TFLN) photonic integrated circuit (PIC) as the external soliton laser cavity. TFLN PICs have recently been shown as a promising platform for a wide range of electro-optic, nonlinear photonic, quantum photonic, and laser applications 37, 38, 39. It exhibits a unique photorefractive (PR) effect that can self-stabilize the optical resonances of a high-Q TFLN microresonator in the red-detuned regime 26. We thus utilize this intriguing feature inside a laser cavity for turn-key mode locking operation as illustrated in Fig. 1a.
The laser comprises three key elements (Fig. 1a,b): InP-based RSOA as the optical gain, a TFLN microring resonator as the Kerr mode-locking element, and a Fabry–Pérot (FP) cavity as the principal laser cavity formed by a broadband Sagnac end mirror on the TFLN PIC side and a reflective end mirror at the facet of the InP RSOA. In this laser, the frequencies of the lasing modes are primarily determined by the FP cavity, while the nonlinear optical effects are dominated by the embedded high-Q microresonator. When the injection current to the RSOA is turned on, multimode lasing waves initiated from the III-V optical gain will pass through the embedded high-Q microresonator, which produces a strong PR effect to blue-shift the optical resonances of the microresonator, resulting in the FP cavity modes located at the red-detuned side of those of the embedded microresonator. At the same time, the strong optical Kerr effect inside the microresonator converts these lasing waves into broadband OFC via the resulting four-wave mixing process. As the produced OFC resides directly in the red-detuned regime of the microresonator ideal for soliton mode locking, the OFC would naturally evolve into mode-locked soliton pulses.
@@
Fig. 1b-d show an example of the fabricated laser device. The high‑Q TFLN microring features an instrinsic optica Q of 2.7 million (See Supplementary Information (SI), Fig. S2a) and a free spectral range (FSR) of 200 GHz, about 77 times of that of the principal FP cavity (FSR of $\sim$2.6 GHz). To characterize the laser performance, the laser output is monitored through the bus waveguides coupled to the microring (designed to be over coupled, with a loaded optical Q of $\sim$1 million). We examined the optical and RF spectra across various operating states by systematically varying the RSOA current while manually optimizing the coupling position to compensate for thermal-induced misalignment between the gain chip and TFLN chip.
@@
As shown in Fig. 1a, the spectrum of the soliton combs broadens considerably with increased RSOA current, reaching a 3-dB optical bandwidth of $\sim$3.4 THz when the current increases to 132 mA. This spectral bandwidth corresponds to a pulse width as small as $\sim$90 fs, significantly shorter than that can be achieved in a conventional SMLL 7, 8. Moreover, the recorded soliton comb spectrum closely matches the spectral characteristics of DKS states produced in a separate passive microring with identical dispersion design and similar coupling conditions, which further validates the soliton nature of these mode-locked comb states. The details are provided in the SI.
@@
To further verify the mode-locked nature of the soliton microcombs emitted from the laser, we examined their temporal characteristics using frequency-resolved optical gating (FROG). As shown in Fig. 2c, FROG traces at 75 mA and 120 mA reveal highly periodic structures in both temporal and frequency domains (see SI for detailed discussion of FROG traces). For example, the FROG trace at 75 mA shows an interference pulse train with a temporal period of $\sim$0.83 ps that corresponds to one-sixth of the fundamental 5-ps round-trip time of the mode-locking microresonator and a frequency spacing of 1.2 THz that corresponds to six times the fundamental 0.2-THz FSR. The FROG trace at 120 mA shows similar behavior while with different temporal period of 0.625 ps and frequency spacing of 1.6 THz. These observations clearly verify the coherent nature of the soliton comb in both time and spectral domains. Moreover, the clean FROG traces indicate that the soliton pulse trains are fully background-free, in stark contrast to DKS that is inevitably accompanied with a strong pump wave background. This can be seen more clearly by comparison with the FROG traces of DKS produced in a passive microresonator, whose details are provided in SI.
@@
When the RSOA current increases to 151 mA, the microcomb spectrum is further broadened with a comb mode spacing of 0.4 THz. FROG characterization reveals a periodic but complex temporal and frequency structure, implying a partially mode-locked soliton state. The complex FROG trace infers a dual-pulse pattern with a period of 2.5 ps that consists of a stable primary pulse and an unstable satellite pulse. The satellite pulse is likely unstable in phase and relative temporal position, manifesting in the FROG trace as pulse peaks residing on a temporally broadened pedestal. The instability of the satellite pulse leads to an elevated level in the detected RF spectrum. The intriguing unstable double-pulsing behavior is reminiscent of the over-pumping conditions or insufficient mode locking (e.g., small reverse bias in SA in conventional SMLLs 42, 43 as well as soliton rain in fiber soliton lasers 44). A detailed investigation of this phenomenon lies beyond the scope of this work and will be explored in the future.
@@
The comb power is measured to be about 1.4 mW on chip with a RSOA driving current of 132 mA. Given that the laser has two output ports (related to two bus waveguides coupled to the microring. Fig. 1), the total output comb power is about 1.5 mW. It corresponds to a wall-plug efficiency of $\sim$1.1% that is on par with conventional SMLLs 7, 8, while here our laser exhibits markedly broader 3-dB optical bandwidth (shorter pulse width) and significantly higher repetition rate. Fig. 2d compares our laser with state-of-the-art SMLLs, which clearly shows the superior performance beyond the reach of conventional SMLLs. The enhanced performance directly arises from the ultrafast, instantaneous soliton mode‑locking mechanism and the effectively unlimited parametric bandwidth supported by our design. The comb power is also on par with those produced with DKS approaches 7, 8, while here our laser requires an overall electric power orders of magnitude lower. Interestingly, we estimate the circulating optical power inside the principal FP cavity to be only around 3 mW, about an order of magnitude lower than the optical power required to excite similar soliton combs under conventional DKS schemes (see SI, Fig. S2). This enhanced efficiency stems from the multi-mode pumping mechanism, where the soliton threshold is inversely proportional to the number of coherent in-phase excitation modes 22. In our device, the broadband laser gain simultaneously supports multiple lasing modes, effectively creating a natural multi-mode pumping for the embedded mode-locking microring resonator (Fig. 1a).
@@
The practical deployment of soliton microcomb sources beyond laboratory settings demands robust, deterministic operation that eliminates the need for delicate tuning procedures. To address this critical requirement, we investigated the turn-key capabilities of our soliton laser, examining both the reliability of soliton self-creation and the underlying temporal dynamics governing the initialization process. We performed multiple on‑off cycling while simultaneously monitoring both the comb power and the heterodyne beatnote with an external reference CW laser, as shown in Fig. 3a. To ensure the beatnote accurately represented the comb state, we selected an off-center comb line (wavelength of 1550 nm) for the 1.2-THz soliton state at 100 mA for heterodyning. As shown in Fig. 3b, the beatnote consistently recurs at an identical frequency and the output optical power returns to the same level across multiple startup and shutdown cycles. This repeatable, “set‑and‑forget” operation demonstrates the reliability of our soliton microcomb laser.
@@
Notably, during each startup ramp, we observe a transient overshoot in comb power and a brief chaotic beatnote before the system settles into the steady soliton comb state. To probe these fast dynamics, we recorded the detailed output power waveform during a current ramp. During the current ramp from 0.5 s to 2 s, lasing commences at a RSOA current of $\sim 50$ mA, followed by several distinct power drop steps representing transitions between cavity lasing states (Fig. 3c). After reaching the 100 mA setpoint, the temporal trace exhibits larger, slower fluctuations before stabilizing at the characteristic power level of the soliton comb. The power variation displays a time constant of about tens to several hundred milliseconds, which is in agreement with our previous measurements of the PR effect on the TFLN platform45. This temporal behavior further confirms the pivotal role of photorefraction in facilitating reliable self‑starting operation.
@@
To further show the detailed lasing dynamics, we finely tune the RSOA current around 100 mA to investigate the 1.2-THz soliton state. As shown in Fig. 3d, below 99 mA, the laser operates in a single-frequency CW state at a wavelength around 1585 nm and with a laser output power of $\sim$1 mW on chip. It is likely because the microring has a stronger external coupling to the bus waveguides at a longer wavelength and thus a lower intracavity loss. As a result, the laser tends to oscillate at the red side of the C-band optical gain when only a single mode is supported. Around this spectral regions, a microring resonance $\omega_{r1}$ coincides with a FP-cavity resonance $\omega_{\rm FP1}$, leading to a minimal loss and single-frequency lasing. At 99 mA, the laser power briefly spikes, then drops sharply before settling at $\sim$0.75 mW, while the optical spectrum abruptly broadens to an equally-spaced comb spectrum centered around 1550 nm, with an ideal $sech^{2}$ envelope characteristic of a soliton comb. With increased RSOA current and lasing power, the PR effect progressively blue-shifts the microresonator such that around a certain different resonance $\omega_{r2}$, the FP cavity mode $\omega_{\rm FP2}$ exhibits appropriate red detuning from $\omega_{r2}$ that supports soliton formation, initiating sudden transition into the mode-locked soliton state with spectrum centered around $\omega_{\rm FP2}$. The observed power dynamics around 99 mA—a transient increase followed by a decline—suggest the system crosses the soliton existence threshold, and then stabilizes into the soliton state. The mode-locked soliton comb state remains stable when the current increases to about 103 mA beyond which the lasing state becomes unstable. We attribute this to excessive detuning—driven by the strengthened PR effect, pushing the laser away from the stable soliton regime. See SI for further discussions, which also provides detailed evolution of optical and RF spectra of the laser output during this transition dynamics.
@@
The soliton state remains very stable as soon as it is turned on. Fig. 3e shows an example for a 1.2-THz soliton state, which exhibits a broad 3-dB bandwidth of 3.9 THz (corresponding to a pulse width of $\sim$80 fs). The state remains stable over the entire recording time period of about 46 minutes. As shown in Fig. 3e-II, it is perturbed only by the mechanical drift of the testing station (as the RSOA gain chip and the TFLN PIC are placed on separate mechanical stages as shown in Fig. 1c). The soliton state is fully recovered after the small mechanical drift is corrected manually (Fig. 3e-III).

### v2

# Chip-scale ultrafast soliton laser

@@ Abstract
Femtosecond laser, owing to their ultrafast time scales and broad frequency bandwidths, have substantially changed fundamental science over the past decades, from chemistry and bio-imaging to quantum physics. Critically, many emerging industrial-scale photonic technologies—such as optical interconnects, AI accelerators, quantum computing, and LiDAR—also stand to benefit from their massive frequency parallelism. However, achieving a femtosecond-scale laser on-chip, constrained by size and system power input, has remained a long-standing challenge. Here, we demonstrate the first on-chip femtosecond laser, enabled by a new mechanism—photorefraction-assisted soliton (PAS) mode-locking. Operating from a simple, low-voltage electrical supply, the laser provides deterministic, turn-key generation of sub-90-fs solitons. Furthermore, it provides electronic reconfigurability of its pulse properties and features an exceptional optical coherence with a 53 Hz intrinsic comb linewidth. This demonstration removes a key barrier to the full integration of chip-scale photonic systems for next-generation sensing, communication, metrology, and computing.
@@ Introduction
Femtosecond laser, which produce femtosecond pulses in the time domain while simultaneously forming broadband coherent optical frequency combs in the spectral domain, have become a foundational technology in modern science1, 2, 3. Their unique capabilities, including ultrafast time scales, massive frequency parallelism, and unparalleled frequency precision, underpin broad cross-discipline applications developed, from molecular chemistry4 and bio-imaging5 to communications6, precision sensing7, 8, and optical clocks9. The impact of these ultrafast broadband lasers extends beyond fundamental science in the laboratory, promising to transform emerging industry-scale photonic technologies that rely on multi-wavelength sources, such as optical interconnects10, 11, autonomous driving12, 5G/6G communication13, 14, and AI accelerators15, 16. Realizing femtosecond-scale broadband combs on-chip, within constrained space and power budgets, is critical for enabling the system parallelism, energy efficiency, and integration required to make these applications practical and mass-producible.
However, realizing femtosecond-pulse source on photonic chips remains fundamentally challenging. Current on-chip mode-locked lasers (MLLs) based on semiconductor platforms (Fig. 1a) offer excellent scalability and, crucially, spontaneous soliton generation from a simple direct-current bias17, 18, 19. However, their pulse performance is fundamentally constrained by the finite carrier relaxation times inherent to saturable absorption (SA) mode-locking, leading to sub-ps scale at best. Alternatively, externally-driven dissipative Kerr solitons (DKS) in high-Q microresonators (Fig. 1b) can achieve femtosecond-scale pulses and high comb line coherence20, 21, and have been applied to many areas22. However, DKS is essentially not a self-contained laser and relies critically on external laser pumping for proper operation. This external-driving scheme introduces fundamental challenges including operational complexity, pump residue, and power inefficiency. Current solutions to these issues generally sacrifice the core benefits of integration by requiring complex and bulky external control elements or sophisticated fiber components for the tuning, stabilization, and synchronization of frequency and time 23, 24, 25, 26, 27, 28, 29, leading to significant system complexity and substantial additional power budget. To date, a fully integrated, power-efficient on-chip light source capable of generating femtosecond-scale pulses has remained elusive.
In this work, we demonstrate the first chip-scale femtosecond laser. This on-chip soliton laser is based on a new mechanism: photorefraction-assisted soliton (PAS) mode-locking. As a result, the laser achieves pulse durations shorter than 90 fs with a repetition rate up to 3 THz. Operating from a simple low-voltage electrical supply, this on-chip laser exhibits deterministic, turn-key soliton generation and outstanding long-term stability. Moreover, the device is electronically reconfigurable, allowing its repetition rate and spectral bandwidth to be tuned via the pump current. The laser also exhibits outstanding long-term stability and a 53-Hz intrinsic comb linewidth, a coherence three orders of magnitude better than the current state-of-the-art for on-chip MLLs.
Figure 1: Concept and implementation of on-chip soliton laser. (a, b) state-of-the-art approaches to chip-scale ultrashort pulse generation. (a) On-chip MLLs. Typically based on III-V platforms, these devices use a reverse-biased semiconductor saturable absorber to achieve self-starting pulse generation directly from an electrical current. However, their pulse duration is fundamentally limited to the sub-picosecond domain by the finite carrier response time of the semiconductor SA. (b) Externally-driven DKS. A high-Q microresonator is pumped by an external continuous-wave (CW) laser to have solitons through four-wave mixing (FWM). While this approach generates femtosecond pulses, the externally-driven parametric process is critically reliant on the external pump laser, introducing challenges such as difficult soliton initiation, low conversion efficiency, pedestal-present soliton, and pump-dependent comb linewidth. (c) Chip-scale femtosecond soliton laser enabled by PAS mode-locking. (c.I) Device architecture. A microring exhibiting a PR effect is embedded within a semiconductor-gain-based FP external cavity. (c.II) Frequency-domain picture of the soliton mode-locking mechanism. (c.III) Time-domain behavior of the mode-locking. (d) Schematic of the hybrid integrated laser architecture. (e,f) Photograph showing the hybrid integration and the device footprint in comparison with a one-cent coin. (g) Micrographs detailing the RSOA-TFLN coupling interface, embedded high-Q microresonator, and broadband Sagnac mirror.
@@
Figure 2: Performance of the soliton microcomb laser. (a) Optical (left) and RF (right) spectra of laser output at different RSOA driving currents. Dashed lines represent $sech^{2}$ fits. Each state was obtained by optimizing the relative position between the RSOA and TFLN chips to compensate for thermal-induced mechanical drift. (b) FROG traces for the 1.2-THz (top), 1.6-THz (middle), and 0.4-THz (bottom) soliton states at RSOA driving currents of 75, 120, and 151 mA, respectively. The center frequency ($\Delta\nu=0$) of the FROG traces is 386.83THz (corresponding wavelength of 775 nm). For a better comparison, the figures also show frequency and time scales normalized, respectively, by the 200-GHz FSR and 5-ps roundtrip time ($T_{0}$) of the mode-locking resonator. The relatively indistinct pulse temporal boundaries in the FROG traces result simply from limited gain bandwidth of the EDFA used to boost the soliton pulses before they are recorded by FROG (Fig. 2d). This artifact is confirmed by performing the same FROG on similar comb states produced via conventional DKS excitation in a passive microresonator. Details are provided in SI. (c) Comparison with advanced on-chip MLLs on pulse duration and repetition rate. All data are listed in Table S1 of SI. The 132 mA state is 90 fs and 2 THz. The 80 fs state is given later in Fig. 3e. (d) Schematic of experimental setup. LDC: laser diode controller; DCU: dispersion compensation unit; PM: power meter; FROG: frequency-resolved optical gating; OSA and ESA: optical and electrical spectrum analyzer.
@@
Our laser integrates an InP‑based reflective semiconductor optical amplifier (RSOA) as the gain medium with a thin-film lithium-niobate (TFLN) photonic integrated circuit (PIC) as the external soliton laser cavity (Fig. 1c.I). TFLN PICs have recently been shown as a promising platform for a wide range of electro-optic, nonlinear photonic, quantum photonic, and laser applications 30, 31, 32. It exhibits a unique photorefractive (PR) effect that can self-stabilize the optical resonances of a high-Q TFLN microresonator in the red-detuned regime 33. We thus utilize this intriguing feature inside a laser cavity for turn-key mode locking operation as illustrated in Fig. 1c.
The laser comprises three key elements (Fig. 1d): InP-based RSOA as the optical gain, a TFLN microring resonator as the Kerr mode-locking element, and a Fabry–Pérot (FP) cavity as the principal laser cavity formed by a broadband Sagnac end mirror on the TFLN PIC side and a reflective end mirror at the facet of the InP RSOA. In this architecture, the FP cavity primarily determines the lasing mode frequencies, while the embedded high-Q microresonator governs the nonlinear optical dynamics.
Upon applying an injection current to the RSOA, multimode lasing waves initiated from the III-V optical gain will pass through the embedded high-Q microresonator, which produces a strong PR effect to blue-shift the optical resonances of the microresonator. This process intrinsically stabilizes the system by forcing the FP cavity modes to operate on the red-detuned side of the microresonator’s resonances (Fig. 1c.II). Simultaneously, the strong optical Kerr effect inside the microresonator converts these lasing waves into a broadband frequency comb via the resulting four-wave mixing process. As the comb is produced in the stable red-detuned regime ideal for soliton formation, it would naturally evolve into a mode-locked soliton pulse train.
The PAS mode-locking can also be seen clearly in the time domain, as illustrated in Fig. 1c.III. Self-phase modulation (SPM) induced by optical pulses dynamically shifts the instantaneous microring resonance (blue line) toward the FP cavity mode (red line). This transient, self-regulating mechanism creates a net-gain window on an ultrafast timescale, favoring femtosecond pulse formation. The PR effect counteracts thermal nonlinearities within the microring mode-locker, which is essential for establishing the stable red-detuned condition required for this ultrafast Kerr modulation. PAS mode-locking can thus be understood as a form of Kerr mode-locking, a mechanism previously proven to be highly effective for realizing femtosecond lasers in free-space and fiber systems1, 34. This work represents the first demonstration of Kerr mode-locking on-chip.
@@
Fig. 1d-f show an example of the fabricated laser device. The high‑Q TFLN microring features an instrinsic optica Q of 2.7 million (See Supplementary Information (SI), Fig. S2a) and a free spectral range (FSR) of 200 GHz, about 77 times of that of the principal FP cavity (FSR of $\sim$2.6 GHz). To characterize the laser performance, the laser output is monitored through the bus waveguides coupled to the microring (designed to be over coupled, with a loaded optical Q of $\sim$1 million). We examined the optical and RF spectra across various operating states by systematically varying the RSOA current while manually optimizing the coupling position to compensate for thermal-induced misalignment between the gain chip and TFLN chip.
@@
As shown in Fig. 2a, the spectrum of the soliton combs broadens considerably with increased RSOA current, reaching a 3-dB optical bandwidth of $\sim$3.4 THz when the current increases to 132 mA. This spectral bandwidth corresponds to a pulse width as small as $\sim$90 fs, significantly shorter than that can be achieved in a conventional on-chip MLL 37, 38. Moreover, the recorded soliton comb spectrum closely matches the spectral characteristics of DKS states produced in a separate passive microring with identical dispersion design and similar coupling conditions, which further validates the soliton nature of these mode-locked comb states. The details are provided in the SI.
@@
To further verify the mode-locked nature of the soliton emitted from the laser, we examined their temporal characteristics using frequency-resolved optical gating (FROG). As shown in Fig. 2b, FROG traces at 75 mA and 120 mA reveal highly periodic structures in both temporal and frequency domains (see SI for detailed discussion of FROG traces). For example, the FROG trace at 75 mA shows an interference pulse train with a temporal period of $\sim$0.83 ps that corresponds to one-sixth of the fundamental 5-ps round-trip time of the mode-locking microresonator and a frequency spacing of 1.2 THz that corresponds to six times the fundamental 0.2-THz FSR. The FROG trace at 120 mA shows similar behavior while with different temporal period of 0.625 ps and frequency spacing of 1.6 THz. These observations clearly verify the coherent nature of the soliton comb in both time and spectral domains. Moreover, the clean FROG traces indicate that the soliton pulse trains are fully background-free, in stark contrast to DKS that is inevitably accompanied with a strong pump wave background. This can be seen more clearly by comparison with the FROG traces of DKS produced in a passive microresonator, whose details are provided in SI.
@@
When the RSOA current increases to 151 mA, the microcomb spectrum is further broadened with a comb mode spacing of 0.4 THz. FROG characterization reveals a periodic but complex temporal and frequency structure, implying a partially mode-locked soliton state. The complex FROG trace infers a dual-pulse pattern with a period of 2.5 ps that consists of a stable primary pulse and an unstable satellite pulse. The satellite pulse is likely unstable in phase and relative temporal position, manifesting in the FROG trace as pulse peaks residing on a temporally broadened pedestal. The instability of the satellite pulse leads to an elevated level in the detected RF spectrum. The intriguing unstable double-pulsing behavior is reminiscent of the over-pumping conditions or insufficient mode locking (e.g., small reverse bias in SA in conventional on-chip MLLs 39, 40 as well as soliton rain in fiber soliton lasers 41). A detailed investigation of this phenomenon lies beyond the scope of this work and will be explored in the future.
@@
The comb power is measured to be about 1.4 mW on chip with a RSOA driving current of 132 mA. Given that the laser has two output ports (related to two bus waveguides coupled to the microring shown in Fig. 1), the total output comb power is about 1.5 mW. It corresponds to a wall-plug efficiency of $\sim$1.1% that is on par with conventional on-chip MLLs 37, 38, while here the soliton laser exhibits markedly broader 3-dB optical bandwidth (shorter pulse width) and significantly higher repetition rate. Fig. 2c compares this laser with state-of-the-art on-chip MLLs, which clearly shows the superior pulse performance beyond the reach of conventional on-chip MLLs. The enhanced performance directly arises from the ultrafast soliton mode‑locking mechanism and the effectively unlimited parametric bandwidth supported by design. The comb power is also on par with those produced with DKS approaches 37, 38, while here the soliton laser requires an overall electric power orders of magnitude lower. Interestingly, we estimate the circulating optical power inside the principal FP cavity to be only around 3 mW, about an order of magnitude lower than the optical power required to excite similar soliton combs under conventional DKS schemes (see SI, Fig. S2). This enhanced efficiency stems from the multi-mode pumping mechanism, where the soliton threshold is inversely proportional to the number of coherent in-phase excitation modes 42. In the device, the broadband laser gain simultaneously supports multiple lasing modes, effectively creating a natural multi-mode self-pumping for the embedded mode-locking microring resonator (Fig. 1).
@@
We investigated the turn-key capabilities of the soliton laser, examining both the reliability of soliton self-creation and the underlying temporal dynamics governing the initialization process. We performed multiple on‑off cycling while simultaneously monitoring both the comb power and the heterodyne beatnote with an external reference CW laser, as shown in Fig. 3a. To ensure the beatnote accurately represented the comb state, we selected an off-center comb line (wavelength of 1550 nm) for the 1.2-THz soliton state at 100 mA for heterodyning. As shown in Fig. 3b, the beatnote consistently recurs at an identical frequency and the output optical power returns to the same level across multiple startup and shutdown cycles. This repeatable, “set‑and‑forget” operation demonstrates the reliability of the soliton microcomb laser.
@@
Notably, during each startup ramp, we observe a transient overshoot in comb power and a brief chaotic beatnote before the system settles into the steady soliton comb state. To probe these fast dynamics, we recorded the detailed output power waveform during a current ramp. During the current ramp from 0.5 s to 2 s, lasing commences at a RSOA current of $\sim 50$ mA, followed by several distinct power drop steps representing transitions between cavity lasing states (Fig. 3c). After reaching the 100 mA setpoint, the temporal trace exhibits larger, slower fluctuations before stabilizing at the characteristic power level of the soliton comb. The power variation displays a time constant of about tens to several hundred milliseconds, which is in agreement with our previous measurements of the PR effect on the TFLN platform43. This temporal behavior further confirms the pivotal role of photorefraction in facilitating reliable self‑starting operation.
@@
To detail the lasing dynamics, we finely investigating the 1.2-THz soliton state at 100 mA (Fig. 3d). Below 99 mA, the laser operates in a single-frequency continuous-wave (CW) state at $\sim$1585 nm, with an on-chip output power of $\sim$1 mW. This long-wavelength operation is likely favored because the microring’s external coupling is stronger at longer wavelengths, resulting in lower intracavity loss at the red side of the C-band gain. At the 99 mA threshold, the laser dynamics change abruptly. The output power briefly spikes, then drops sharply before settling at $\sim$0.75 mW. Simultaneously, the optical spectrum transitions from a single CW line to a broadband, equally-spaced comb centered at $\sim$1560 nm, which exhibits the ideal $sech^{2}$ envelope characteristic of a soliton. This transition is driven by the PR effect: as the RSOA current and intracavity power increase, the PR effect progressively blue-shifts the microresonator resonances. At 99 mA, some FP cavity modes achieves the optimal red-detuning relative to microresonator resonances, enabling and initiating the sudden transition into the mode-locked soliton state. The observed power dynamics—a transient spike followed by a stable drop—are a classic signature of the system crossing the soliton existence threshold. This mode-locked soliton comb state remains stable as the current increases to $\sim$103 mA. Beyond this point, the lasing becomes unstable. We attribute this to excessive detuning driven by the strengthened PR effect, pushing the laser away from the stable soliton regime. The detailed evolution of the optical and RF spectra during this transition is provided in the SI.
@@
The soliton state remains very stable as soon as it is turned on. Fig. 3e shows an example for a 1.2-THz soliton state, which exhibits a broad 3-dB bandwidth of 3.9 THz (corresponding to a pulse width of $\sim$80 fs). The state remains stable over the entire recording time period of about 46 minutes. As shown in Fig. 3e-II, it is perturbed only by the mechanical drift of the testing station (as the RSOA gain chip and the TFLN PIC are placed on separate mechanical stages as shown in Fig. 1d). The soliton state is fully recovered after the small mechanical drift is corrected manually (Fig. 3e-III).